OpenStudy (anonymous):

EASY!! AWARD WILL BE GIVIN! Question 21 (Multiple Choice Worth 3 points) [03.06] Which is the most profitable investment for a cell phone company earning $10 per plan contract? opening a new store location with $200 daily operating cost, generating 30 new plans per day redesigning web site at $50 daily operating cost, generating 10 new plans per day creating television commercial at $25 daily operating cost, generating 15 new plans per day offering free $10 gift for every new plan, generating 40 new plans per day

OpenStudy (anonymous):

1. 300-200=100 2.100-50=50 3.150-25=125 4.400-400=0 So plan 3. Hope I understood the question correct.

OpenStudy (anonymous):

thts a cat:(

OpenStudy (primeralph):

Suddenly, a dancing CAT appears _ヽ   \\ Λ_Λ    \( ˇωˇ)      > ⌒ヽ    /   へ\    /  / \\    レ ノ   ヽ_つ   / /   / /|  ( (ヽ  | |、\  | 丿 \ ⌒)  | |  ) / `ノ )  Lノ (_/.